# Generation of large-scale magnetic fields upstream of gamma-ray burst afterglow shocks

### v1

@@ 3 Case study: $\alpha=2,\;t_{f}\omega_{pb}=10^{4}$
We begin by describing the salient features of a reference simulation with $\alpha=2$ and $t_{f}\omega_{pb}=10^{4}$; this case demonstrates the various stages of magnetic field evolution that we see in each of our simulations. As we inject electron-positron pairs over the duration of the simulation, the magnetic field evolves through three general stages, as summarized in Figure 1. Figure 1a shows the complete time evolution of the energy densities of the magnetic and electric fields (blue and orange curves, respectively); we normalize these energy densities to the bulk kinetic energy of the injected pairs. We define the box-averaged magnetic energy density to be
$\varepsilon_{B}\equiv\frac{\left<B_{z}^{2}\right>/8\pi}{(\gamma_{b}-1)m_{e}n_{b}c^{2}},$ (3)
@@
where $B_{z}$ is the $z$-component of the magnetic field (the only non-zero component in our 2D unmagnetized plasma), $\gamma_{b}$ is the bulk Lorentz factor of the injected pairs, and $n_{b}$ is the total beam number density at the end of injection (in this case, $\gamma_{b}=1.5$ and $n_{b}$ is twice the background number density). Similarly, we define the box-averaged electric energy densities to be
$\varepsilon_{E_{i}}\equiv\frac{\left<E_{i}^{2}\right>/8\pi}{(\gamma_{b}-1)m_{e}n_{b}c^{2}},$ (4)
@@
Alongside the energy densities in Figure 1a, we also show the time evolution of $k_{y}\right>$ – the average of the $y$-component of the magnetic field wavenumber (i.e., the magnetic wavenumber transverse to the beam) – plotted in green and measured on the right vertical axis in units of $\left(c/\omega_{pb}\right)^{-1}$; we start plotting $k_{y}\right>$ at $t_{f}\omega_{pb}\sim 1100$, when the magnetic energy overtakes the electric energy. $k_{y}\right>$ is derived from the 1D transverse power spectrum of the magnetic field, $P_{B}(k_{y})$:
@@
We obtain $P_{B}(k_{y})$ by averaging the 2D magnetic power spectrum over $k_{x}$. We place a finite upper bound, $k_{\rm cut}$, on the integrals in Equation (5) to avoid contaminating our measurements with the high-$k_{y}$ end of the magnetic spectrum, which is affected by numerical shot noise; we choose $k_{\rm cut}$ to be the location where $k_{y}P_{B}(k_{y})$ reaches a minimum between the peak of $k_{y}P_{B}(k_{y})$ and $k_{y}c/\omega_{pb}=1$. We find that placing similar bounds on $k_{x}$ negligibly affects our results, so the quantities computed in this paper use the full range of $k_{x}$.
@@
Figure 1: Time evolution and a few representative snapshots from our reference simulation with $\alpha=2$ and $t_{f}\omega_{pb}=10^{4}$. Panel (a) shows the full time evolution of the box-averaged magnetic energy density ($\varepsilon_{B}$, solid blue curve) and electric energy densities ($\varepsilon_{E_{x}}$ and $\varepsilon_{E_{y}}$, dashed orange and dotted orange curves, respectively), as well as the average transverse magnetic wavenumber ($k_{y}\right>c/\omega_{pb}$, green curve, measured on the right vertical axis for times later than $t\omega_{pb}\sim 1100$, when the magnetic energy overtakes the electric energy); the arrow at $k_{y}\right>c/\omega_{pb}\sim 10^{-1}$ on the right vertical axis indicates the characteristic wavenumber $2\pi/(c/\omega_{pi})$. The four vertical grayscale lines mark the times corresponding to the snapshots below, with each line corresponding to a different row. Panels (b) and (c) show the 2D spatial profile and 2D power spectrum of $E_{x}$, respectively, at $t\omega_{pb}\sim 770$, the peak of the electrostatic phase. Panels (d) and (e) show the 2D spatial profile and the 2D power spectrum of $B_{z}$, respectively, at $t\omega_{pb}\sim 1100$, when the magnetic energy overtakes the electric energy. Panels (f) and (g) again show the spatial profile and power spectrum of $B_{z}$, but at $t\omega_{pb}\sim 1850$, when the filamentation phase saturates. Finally, panels (h) and (i) show the spatial profile and power spectrum of $B_{z}$ at an arbitrarily late time, $t\omega_{pb}\sim 8280$, deep into the secular phase. In panels (d), (f), and (h), the color bar is rescaled to $\left|B_{z}\right|^{0.2}\times{\rm sign}(B_{z})$ to display a wider dynamic range.
At early times (i.e., $t\omega_{pb}\lesssim 1100$), the total energy of the system is dominated by the electric field, with $\varepsilon_{E_{x}}>\varepsilon_{E_{y}}\gg\varepsilon_{B}$. This early exponential growth of the electric energy is driven by the development of both the electrostatic two-stream instability and the quasi-electrostatic oblique instability (Bret et al., 2010a), as evidenced by the spatial profile of $E_{x}$ (normalized by $\sqrt{8\pi\left(\gamma_{b}-1\right)m_{e}n_{b}c^{2}}$) and its associated 2D power spectrum in Figures 1b and 1c, respectively. The oscillations with $x$-aligned (i.e., beam-aligned) wavevectors in Figure 1b and the power at $(k_{x}c/\omega_{pb}\sim\pm 1,\,k_{y}c/\omega_{pb}=0)$ in Figure 1c are indicative of two-stream modes, while the oblique oscillations in Figure 1b and the power at $(k_{x}c/\omega_{pb}\sim\pm 1,\,k_{y}c/\omega_{pb}\sim\pm 1)$ in Figure 1c are indicative of oblique modes. The subdominant – but still exponential – growth of the magnetic energy is due to the electromagnetic component of the oblique mode.
@@
Figure 2: Time evolution of a few representative wavemodes plotted alongside $\varepsilon_{B}$ (blue curve; arbitrarily rescaled) for our $\alpha=2$, $t_{f}\omega_{pb}=10^{4}$ reference simulation. $k_{y}c/\omega_{pb}\sim 0.408$ (orange curve) and $k_{y}c/\omega_{pb}\sim 0.387$ (green curve) – the two modes that carry the most magnetic power at the saturation of the electrostatic phase, $t_{f}\omega_{pb}\sim 850$ (vertical light gray line) – are quickly overtaken by the much larger-scale modes $k_{y}c/\omega_{pb}\sim 0.052$ (pink curve) and $k_{y}c/\omega_{pb}\sim 0.042$ (brown curve) – the two modes that carry the most power at the saturation of the filamentation phase, $t\omega_{pb}\sim 1850$ (vertical dark gray line).
@@
After the two-stream and oblique modes saturate around $t\omega_{pb}\sim 850$, the magnetic energy enters a second stage of exponential growth from $1000\lesssim t\omega_{pb}\lesssim 1850$, with $\varepsilon_{B}$ quickly overtaking $\varepsilon_{E_{x}}$ and $\varepsilon_{E_{y}}$ by $t\omega_{pb}\sim 1100$. Figures 1d and 1e show, respectively, the spatial profile of $B_{z}$ (normalized by $\sqrt{8\pi\left(\gamma_{b}-1\right)m_{e}n_{b}c^{2}}$) and the 2D power spectrum of $B_{z}$ at the time when the magnetic energy surpasses the electric energy. The thin structures stretched along the $x$-direction in Figure 1d and the dumbbell shape oriented along $k_{x}c/\omega_{pb}=0$ in Figure 1e are characteristic of the filamentation instability. Figure 2 plots the power carried by a few individual transverse wavemodes over time, revealing that the envelope of $\varepsilon_{B}$ as plotted in Figure 1a is shaped by progressively larger-scale modes as the growth proceeds.
@@
As the phase of filamentation growth proceeds from $1100\lesssim t\omega_{pb}\lesssim 1850$, the average transverse scale of the magnetic field increases by more than a factor of two, as shown by the drop in $k_{y}\right>c/\omega_{pb}$ in Figure 1a. This increase in scale can also be seen in Figures 1f and 1g, which depict the spatial profile and 2D power spectrum of $B_{z}$ at $t\omega_{pb}\sim 1850$, where the filamentation growth saturates; Figures 1f and 1g can be directly compared to the earlier snapshots in Figures 1d and 1e. Interestingly, the formerly ordered filamentary structures of Figure 1d appear to be turbulently mixed as the filamentation growth saturates (Figure 1f), but this order is restored at later times, albeit on much larger scales (Figure 1h).
@@
Figure 3: Time evolution of the box-averaged $x$-velocities ($v_{x}/c$, panel (a)) and temperatures ($k_{B}T/m_{e}c^{2}$, panel (c)) for the beam electrons (blue), beam positrons (orange), background electrons (green), and background protons (pink) in our $\alpha=2$, $t_{f}\omega_{pb}=10^{4}$ reference simulation. Panel (b) zooms in on the proton velocity curve from panel (a). In each panel, the light gray vertical line marks the time of saturation of the electrostatic phase and the dark gray vertical line marks the time of saturation of the filamentation phase.
@@
After the saturation of the filamentation phase at $t\omega_{pb}\sim 1850$, the magnetic energy dips slightly as magnetic filaments break up and form cavities. However, these large-scale filaments appear again at later times – as illustrated by the late-time growth of the small-$k_{y}$ modes in Figure 2 – leading to a final phase of slow, secular growth of the magnetic field from $t\omega_{pb}\sim 4000$ until the end of the simulation at $t\omega_{pb}=10^{4}$.
@@
As the filamentation phase saturates at $t\omega_{pb}\sim 1850$, momentum is rapidly transferred from the beam to the background plasma, accelerating and heating the background electrons and protons; this is illustrated in Figure 3, which shows the velocities and temperatures of each particle species over time in our reference simulation. The background electrons and protons – which had largely been stationary at earlier times – start picking up small bulk $x$-velocities at $t_{f}\omega_{pb}\sim 1600$, but decouple due to their mass difference. This background mass difference also causes the beam electrons to moderately decouple from the beam positrons; such symmetry breaking is known to drive the growth of low-density, high-magnetic-energy plasma cavities, which we see in Figure 1f (Peterson et al., 2021, 2022; Grošelj et al., 2024). Meanwhile, the relative drift between the protons and the leptons triggers a secondary filamentation instability that drives the observed late-time magnetic field growth, as evidenced by the formation of proton filaments that were absent at the peak of the first filamentation phase (contrast the left and right columns of Figure 8). This effect is more pronounced at large $\alpha$, where the properties of the background electrons become nearly the same as those of the beam species following saturation of the filamentation phase; as we show in Section 4, increasing $\alpha$ (at fixed $t_{f}$) increases both the secular phase growth rate and the density contrast of the proton filaments.
@@
In summary, the evolution of the magnetic field proceeds through three phases: first, a primarily electrostatic phase dominated by two-stream and oblique modes ($t\omega_{pb}\lesssim 1100$); second, an exponential electromagnetic phase driven by filamentation modes enhanced by continuous pair injection ($1100\lesssim t\omega_{pb}\lesssim 1850$), during which the scale of the magnetic field increases by a factor of more than two; and third, a phase of slow, secular growth driven by the relative drift between the leptons and the protons ($t\omega_{pb}\gtrsim 4000$), at the end of which the scale of the magnetic field is roughly four times larger than at the start of the filamentation phase. In the next section, we show that these phases are generic across the range of $\alpha$ and $t_{f}$ that we simulate.
4 Comparison across $\alpha$ and $t_{f}$

### v2

@@ 3 Case study: $\alpha=2,\;t_{\rm f}\omega_{\rm pb}=10^{4}$
We begin by describing the salient features of a reference simulation with $\alpha=2$ and $t_{\rm f}\omega_{\rm pb}=10^{4}$; this case demonstrates the various stages of magnetic field evolution that we see in each of our simulations. As we inject electron-positron pairs over the duration of the simulation, the magnetic field evolves through three general stages, as summarized in Figure 1. Figure 1a shows the complete time evolution of the energy densities of the magnetic and electric fields (blue and orange curves, respectively); we normalize these energy densities to the bulk kinetic energy of the injected pairs. We define the box-averaged magnetic energy density to be
$\varepsilon_{B}\equiv\frac{\left<B_{z}^{2}\right>/8\pi}{(\gamma_{\rm b}-1)m_{\rm e}n_{\rm b}c^{2}},$ (3)
@@
where $B_{z}$ is the $z$-component of the magnetic field (the only non-zero component in our 2D unmagnetized plasma), $\gamma_{\rm b}$ is the bulk Lorentz factor of the injected pairs, and $n_{\rm b}$ is the total beam number density at the end of injection (in this case, $\gamma_{\rm b}=1.5$ and $n_{\rm b}$ is twice the background number density). Similarly, we define the box-averaged electric energy densities to be
$\varepsilon_{E_{i}}\equiv\frac{\left<E_{i}^{2}\right>/8\pi}{(\gamma_{\rm b}-1)m_{\rm e}n_{\rm b}c^{2}},$ (4)
@@
Alongside the energy densities in Figure 1a, we also show the time evolution of $k_{y}\right>$ – the average of the $y$-component of the magnetic field wavenumber (i.e., the magnetic wavenumber transverse to the beam) – plotted in green and measured on the right vertical axis in units of $\left(c/\omega_{\rm pb}\right)^{-1}$; we start plotting $k_{y}\right>$ at $t_{\rm f}\omega_{\rm pb}\sim 1100$, when the magnetic energy overtakes the electric energy. $k_{y}\right>$ is derived from the 1D transverse power spectrum of the magnetic field, $P_{B}(k_{y})$:
@@
We obtain $P_{B}(k_{y})$ by averaging the 2D magnetic power spectrum over $k_{x}$. We place a finite upper bound, $k_{\rm cut}$, on the integrals in Equation (5) to avoid contaminating our measurements with the high-$k_{y}$ end of the magnetic spectrum, which is affected by numerical shot noise; we choose $k_{\rm cut}$ to be the location where $k_{y}P_{B}(k_{y})$ reaches a minimum between the peak of $k_{y}P_{B}(k_{y})$ and $k_{y}c/\omega_{\rm pb}=1$. We find that placing similar bounds on $k_{x}$ negligibly affects our results, so the quantities computed in this paper use the full range of $k_{x}$. [^2] Alternatively, we could quantify the dominant magnetic wavenumber using the “integral-scale” definition (Equation (D1)), which should be more robust against high-$k_{y}$ noise. In Appendix D, we show that our results change negligibly if we use the integral-scale wavenumber instead of $k_{y}\right>$.
@@
Figure 1: Time evolution and a few representative snapshots from our reference simulation with $\alpha=2$ and $t_{\rm f}\omega_{\rm pb}=10^{4}$. Panel (a) shows the full time evolution of the box-averaged magnetic energy density ($\varepsilon_{B}$, solid blue curve) and electric energy densities ($\varepsilon_{E_{x}}$ and $\varepsilon_{E_{y}}$, dashed orange and dotted orange curves, respectively), as well as the average transverse magnetic wavenumber ($k_{y}\right>c/\omega_{\rm pb}$, green curve, measured on the right vertical axis for times later than $t\omega_{\rm pb}\sim 1100$, when the magnetic energy overtakes the electric energy); the arrow at $k_{y}\right>c/\omega_{\rm pb}\sim 10^{-1}$ on the right vertical axis indicates the characteristic wavenumber $2\pi/(c/\omega_{\rm pi})$. The four vertical grayscale lines mark the times corresponding to the snapshots below, with each line corresponding to a different row. Panels (b) and (c) show the 2D spatial profile and 2D power spectrum of $E_{x}$, respectively, at $t\omega_{\rm pb}\sim 770$, the peak of the electrostatic phase. Panels (d) and (e) show the 2D spatial profile and the 2D power spectrum of $B_{z}$, respectively, at $t\omega_{\rm pb}\sim 1100$, when the magnetic energy overtakes the electric energy. Panels (f) and (g) again show the spatial profile and power spectrum of $B_{z}$, but at $t\omega_{\rm pb}\sim 1850$, when the filamentation phase saturates. Finally, panels (h) and (i) show the spatial profile and power spectrum of $B_{z}$ at an arbitrarily late time, $t\omega_{\rm pb}\sim 8280$, deep into the secular phase. In panels (d), (f), and (h), the color bar is rescaled to $\left|B_{z}\right|^{0.2}\times{\rm sign}(B_{z})$ to display a wider dynamic range.
At early times (i.e., $t\omega_{\rm pb}\lesssim 1100$), the total energy of the system is dominated by the electric field, with $\varepsilon_{E_{x}}>\varepsilon_{E_{y}}\gg\varepsilon_{B}$. This early exponential growth of the electric energy is driven by the development of both the electrostatic two-stream instability and the quasi-electrostatic oblique instability (Bret et al., 2010), as evidenced by the spatial profile of $E_{x}$ (normalized by $\sqrt{8\pi\left(\gamma_{\rm b}-1\right)m_{\rm e}n_{\rm b}c^{2}}$) and its associated 2D power spectrum in Figures 1b and 1c, respectively. The oscillations with $x$-aligned (i.e., beam-aligned) wavevectors in Figure 1b and the power at $(k_{x}c/\omega_{\rm pb}\sim\pm 1,\,k_{y}c/\omega_{\rm pb}=0)$ in Figure 1c are indicative of two-stream modes, while the oblique oscillations in Figure 1b and the power at $(k_{x}c/\omega_{\rm pb}\sim\pm 1,\,k_{y}c/\omega_{\rm pb}\sim\pm 1)$ in Figure 1c are indicative of oblique modes. The subdominant – but still exponential – growth of the magnetic energy is due to the electromagnetic component of the oblique mode.
@@
Figure 2: Time evolution of a few representative wavemodes plotted alongside $\varepsilon_{B}$ (blue curve; arbitrarily rescaled) for our $\alpha=2$, $t_{\rm f}\omega_{\rm pb}=10^{4}$ reference simulation. $k_{y}c/\omega_{\rm pb}\sim 0.408$ (orange curve) and $k_{y}c/\omega_{\rm pb}\sim 0.387$ (green curve) – the two modes that carry the most magnetic power at the saturation of the electrostatic phase, $t_{\rm f}\omega_{\rm pb}\sim 850$ (vertical light gray line) – are quickly overtaken by the much larger-scale modes $k_{y}c/\omega_{\rm pb}\sim 0.052$ (pink curve) and $k_{y}c/\omega_{\rm pb}\sim 0.042$ (brown curve) – the two modes that carry the most power at the saturation of the filamentation phase, $t\omega_{\rm pb}\sim 1850$ (vertical dark gray line).
@@
After the two-stream and oblique modes saturate around $t\omega_{\rm pb}\sim 850$, the magnetic energy enters a second stage of exponential growth from $1000\lesssim t\omega_{\rm pb}\lesssim 1850$, with $\varepsilon_{B}$ quickly overtaking $\varepsilon_{E_{x}}$ and $\varepsilon_{E_{y}}$ by $t\omega_{\rm pb}\sim 1100$. Figures 1d and 1e show, respectively, the spatial profile of $B_{z}$ (normalized by $\sqrt{8\pi\left(\gamma_{\rm b}-1\right)m_{\rm e}n_{\rm b}c^{2}}$) and the 2D power spectrum of $B_{z}$ at the time when the magnetic energy surpasses the electric energy. The thin structures stretched along the $x$-direction in Figure 1d and the dumbbell shape oriented along $k_{x}c/\omega_{\rm pb}=0$ in Figure 1e are characteristic of the filamentation instability. Figure 2 plots the power carried by a few individual transverse wavemodes over time, revealing that the envelope of $\varepsilon_{B}$ as plotted in Figure 1a is shaped by progressively larger-scale modes as the growth proceeds.
@@
As the phase of filamentation growth proceeds from $1100\lesssim t\omega_{\rm pb}\lesssim 1850$, the average transverse scale of the magnetic field increases by more than a factor of two, as shown by the drop in $k_{y}\right>c/\omega_{\rm pb}$ in Figure 1a. This increase in scale can also be seen in Figures 1f and 1g, which depict the spatial profile and 2D power spectrum of $B_{z}$ at $t\omega_{\rm pb}\sim 1850$, where the filamentation growth saturates; Figures 1f and 1g can be directly compared to the earlier snapshots in Figures 1d and 1e. Interestingly, the formerly ordered filamentary structures of Figure 1d appear to be turbulently mixed as the filamentation growth saturates (Figure 1f), but this order is restored at later times, albeit on much larger scales (Figure 1h).
@@
Figure 3: Time evolution of the box-averaged $x$-velocities ($v_{x}/c$, panel (a)) and temperatures ($k_{\rm B}T/m_{\rm e}c^{2}$, panel (c)) for the beam electrons (blue), beam positrons (orange), background electrons (green), and background protons (pink) in our $\alpha=2$, $t_{\rm f}\omega_{\rm pb}=10^{4}$ reference simulation. Panel (b) zooms in on the proton velocity curve from panel (a). In each panel, the light gray vertical line marks the time of saturation of the electrostatic phase and the dark gray vertical line marks the time of saturation of the filamentation phase.
@@
After the saturation of the filamentation phase at $t\omega_{\rm pb}\sim 1850$, the magnetic energy dips slightly as magnetic filaments break up and form cavities. However, these large-scale filaments appear again at later times – as illustrated by the late-time growth of the small-$k_{y}$ modes in Figure 2 – leading to a final phase of slow, secular growth of the magnetic field from $t\omega_{\rm pb}\sim 4000$ until the end of the simulation at $t\omega_{\rm pb}=10^{4}$.
@@
As the filamentation phase saturates at $t\omega_{\rm pb}\sim 1850$, momentum is rapidly transferred from the beam to the background plasma, accelerating and heating the background electrons and protons; this is illustrated in Figure 3, which shows the velocities and temperatures of each particle species over time in our reference simulation. The background electrons and protons – which had largely been stationary at earlier times – start picking up small bulk $x$-velocities at $t_{\rm f}\omega_{\rm pb}\sim 1600$, but decouple due to their mass difference. This background mass difference also causes the beam electrons to moderately decouple from the beam positrons; such symmetry breaking is known to drive the growth of low-density, high-magnetic-energy plasma cavities, which we see in Figure 1f (Peterson et al., 2021, 2022; Grošelj et al., 2024). Meanwhile, the relative drift between the protons and the leptons triggers a secondary filamentation instability that drives the observed late-time magnetic field growth, as evidenced by the formation of proton filaments that were absent at the peak of the first filamentation phase (see Section 4.1). This effect is more pronounced at large $\alpha$, where the properties of the background electrons become nearly the same as those of the beam species following saturation of the filamentation phase; as we show in Section 4.1, increasing $\alpha$ (at fixed $t_{\rm f}$) increases both the secular phase growth rate and the density contrast of the proton filaments.
@@
In summary, the evolution of the magnetic field proceeds through three phases: first, a primarily electrostatic phase dominated by two-stream and oblique modes ($t\omega_{\rm pb}\lesssim 1100$); second, an exponential electromagnetic phase driven by filamentation modes enhanced by continuous pair injection ($1100\lesssim t\omega_{\rm pb}\lesssim 1850$), during which the scale of the magnetic field increases by a factor of more than two; and third, a phase of slow, secular growth driven by the relative drift between the leptons and the protons ($t\omega_{\rm pb}\gtrsim 4000$), at the end of which the scale of the magnetic field is roughly four times larger than at the start of the filamentation phase. In the next section, we show that these phases are generic across the range of $\alpha$ and $t_{\rm f}$ that we simulate.
4 Comparison across $\alpha$ and $t_{\rm f}$